# Structure-Enhanced Protein Instruction Tuning: Towards General-Purpose Protein Understanding

### v2

@@ Appendix A Construction of Protein Instruction Dataset
Currently, the widely recognized protein-text paired databases in the protein domain mainly include Swiss-Prot, TrEMBL (Bairoch & Apweiler, 1997), and RCSB PDB (RCSB, 2024), with their specific contents detailed in Table 6. Since most of the text content in TrEMBL comes from automatic annotation methods, to eliminate the impact of its unreliability on the main experiments, unless otherwise specified, this paper only uses protein-text information from Swiss-Prot and RCSB PDB databases. The corresponding structures are from AlphafoldDB (Varadi et al., 2022) and experimentally determined structures in RCSB PDB. The statistical information about the complete protein instruction dataset is shown in Table ,7 and Figure 4.
Table 6: Protein-text paired database.
@@
Table 7: Statistical information about the protein instruction dataset.
@@
Open-ended Generation Swiss-Prot Specific property or function 2,529,006 11,911
RCSB PDB Protein caption 143,254 2,500
Closed-set Answer RCSB PDB Specific property or function 1,264,286 22,500
GO-BP Biological process 622,935 24,162
GO-MF Molecular function 404,629 5,891
GO-CC Cellular component 334,032 6,735
EC Enzymatic catalytic activity 176,942 2,278
# All Instructions 5,475,084 75,977
@@
Figure 4: Distribution of protein sequence length and conversation length in the protein instruction dataset (training set).
@@
    • What is the primary function of <protein>?
    • What is the main function of <protein>?
    • What is the function of <protein>?
    • Explain the function of <protein>.
    • What is the characteristic function associated with the protein <protein>?
    • Can you define the function profile of the <protein>?
    • Give me the function caption of <protein>.
      Example Response
@@
    • Which protein family does <protein> belong to?
    • What is the protein family of <protein>?
    • What is the closest related protein family for <protein>?
    • Can you identify the family or group that <protein> belongs to?
    • To which protein family <protein> is classified?
    • Which protein class does <protein> fall into?
      Example Response
@@
    • Where is <protein> located in the cell?
    • Can you specify the subcellular location of <protein>?
    • What is the subcellular location of <protein>?
    • Could you describe the subcellular location of <protein>?
    • What are the primary subcellular regions where <protein> is detected?
      Example Response
@@
    • Description the effects of environmental factors of <protein>’s expression.
    • What are the environmental factors that induce the expression of <protein>?
    • What environmental factors causes the upregulation of <protein>?
    • What are the environmental factors that lead to the upregulation of <protein>?
      Example Response
@@
    • Which GO molecular function terms have <protein> been assigned to?
    • What molecular function is associated with <protein>?
    • Which GO terms outline the functional capabilities of <protein>?
    • What are the molecular functions of <protein>?
      Example Response
@@
    • Which GO biological process terms have <protein> been assigned to?
    • What biological process is associated with <protein>?
    • Which GO terms outline the biological processes of <protein>?
    • What biological processes is <protein> involved in, based on gene ontology annotations?
    • What are the biological processes of <protein>?
      Example Response
@@
    • Which GO cellular component terms have <protein> been assigned to?
    • What cellular component is associated with <protein>?
    • Which GO terms outline the cellular components of <protein>?
    • What cellular components is <protein> involved in, based on gene ontology annotations?
    • What are the cellular components of <protein>?
      Example Response
@@
    • At which specific developmental stages is <protein> expressed?
    • What are the developmental stages where <protein> is expressed?
    • What are the developmental stages where <protein> is detected?
    • What are the developmental stages where <protein> is found?
      Example Response
@@
    • Can you identify and list all the motifs that are predicted to be present in <protein>?
    • What are the short sequence motifs that are predicted to be present in <protein>?
    • What are the short sequence motifs that are present in <protein>?
    • What are the short sequence motifs that are found in <protein>?
      Example Response
@@
    • In which tissues is the expression of <protein> absent?
    • Describe the tissue-specific expression pattern of <protein>?
    • What is the tissue-specific expression pattern of <protein>?
    • What are the tissues where <protein> is expressed?
      Example Response
@@
    • Describe the activity regulatory mechanism of <protein> associated enzymes, transporters, microbial transcription factors.
    • What is the activity regulatory mechanism of <protein>?
    • Tell me about the activity regulatory mechanism of <protein>.
      Example Response
@@
    • What is the role of <protein> in the metabolic pathway?
    • Which metabolic pathway does <protein> associate with?
    • What is the metabolic pathway that <protein> is involved in?
      Example Response
@@
    • Tell me about this protein <protein>.
    • Give me some information about <protein>.
    • Give me the abstract of <protein>.
    • Give me a comprehensive description of <protein>.
    • Tell me about <protein>.
      Example Response
@@
    • Does this protein contain non-polymer entities, <protein>?
    • Does this protein contain polymer entities, <protein>?
    • Does this protein contain DNA polymer entities, <protein>?
    • Does this protein contain RNA polymer entities, <protein>?
    • Does this protein contain solvent entities, <protein>?
    • Does this protein contain branched entities, <protein>?
    • Does this protein have unmodeled polymer monomers, <protein>?
    • Does this protein have hybrid nucleic acid polymer entities, <protein>?
    • Does this protein have cis-peptide linkages, <protein>?
      Example Response
@@
    • Does <protein> associate with enzyme classification "{EC}"?
    • Does EC term "{EC}" outline the enzyme classifications of <protein>?
    • Is <protein> involved in enzyme classification "{EC}"?
      Example Response
@@
    • Does <protein> associate with biological process "{GO}"?
    • Does GO term "{GO}" outline the biological processes of <protein>?
    • Is <protein> involved in biological process "GO"?
      Example Response
@@
    • Does <protein> associate with cellular component "{GO}"?
    • Does GO term "{GO}" outline the cellular components of <protein>?
    • Is <protein> involved in cellular component "{GO}"?
      Example Response
@@
    • Does <protein> associate with molecular function "{GO}"?
    • Does GO term "{GO}" outline the functional capabilities of <protein>?
    • Does <pro tein> have molecular function "{GO}"?
      Example Response
@@
Appendix B Supplement to the Experiments
B.1 Implementation Details
As depicted in Figure 2, we extensively utilize group learning rates throughout the entire training pipeline of SEPIT. We tend to assign higher learning rates to randomly initialized parameters, while opting for lower learning rates for pre-trained parameters in order to mitigate forgetting, setting the ratio between lower and higher learning rates at 0.1. In Stage 0, we actually employ ESM2-650M (Lin et al., 2023) as the pLM and PubMedBert (Gu et al., 2021) as the text encoder to better encode biomedical text. We set the number of Gaussian Basis Kernel to 128. In Stage 1, we choose the representation from the penultimate layer of the protein encoder as input to LLMs to minimize the discrepancy between pre-training tasks and the current task. For the LLMs, we opt for TinyLlama-1.1B (Zhang et al., 2024). In Stage 2, we continue most of the settings from Stage 1, while setting the number of experts to 4, with Top-1 expert being activated at a time. At this stage, our protein encoder is frozen. Regarding the hyper-parameter settings for SEPIT-TinyLlama-MoEs, we set higher learning rate to $5e^{-5}$ and trained for 5 epochs in Stage 0. In Stage 1, we set higher learning rate to $2e^{-5}$ and trained for 1 epoch. In Stage 2, we set higher learning rate to $5e^{-5}$ and trained for 1 epoch. For all stages, we trained on 32 Tesla V100 GPUs, with a batch size per GPU set to 4, employing a warm-up and linear decay learning rate scheduler, and set the warm-up ratio to 0.06. In all experiments, we employe AMP, Zero Optimizer (Rasley et al., 2020) based on AdamW (Loshchilov & Hutter, 2019) and gradient checkpointing. For the hyper-parameter settings of other models, see Table 9. For all MoE models, we implement them based on DeepSpeed-MoE, employing expert parallelism during the training process and setting the expert parallel size to 4.
@@ B.1 Implementation Details
Table 8: Tokens consumed by API-based models.
@@
Table 9: Hyper-parameters of all models.
@@
Table 10: Training cost of all models.
@@
Table 11: Inference cost of all models.
@@
B.2 Metric Explanation
BLEU score:
The BLEU (Bilingual Evaluation Understudy) score (Sutskever et al., 2014) is a metric used to evaluate the quality of machine-translated text against human-translated reference texts, which is calculated using n-gram precision. The general formula for calculating BLEU score is as follows, where $\mathbf{BP}$ penalize overly short translations:
$p_{n}=\frac{\sum_{C\in\{\text{ Candidates }\}}\sum_{\text{n-gram }\in C}\text{Count }_{\text{clip }}(\text{ n-gram })}{\sum_{C^{\prime}\in\{\text{ Candidates }\}}\sum_{\text{ n-gram }^{\prime}\in C^{\prime}}\operatorname{Count}\left(\text{ n-gram }^{\prime}\right)},$ (16)
$\mathbf{BP}=\begin{cases}1&\text{if }c>r\\ e^{(1-r/c)}&\text{if }c\leq r\end{cases},$ (17)
$\mathbf{BLEU}=\mathbf{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),$ (18)
$\log\mathbf{BLEU}=\min(1-\frac{r}{c},0)+\sum_{n=1}^{N}w_{n}\log p_{n}.$ (19)
ROUGE-N score:
@@ B.2 Metric Explanation
$\mathbf{ROUGE-N}=\frac{\sum_{C\in\{\text{ Candidates }\}}\sum_{\text{ n-gram }\in C}\operatorname{Count}_{\text{match}}(\text{ n-gram })}{\sum_{C^{\prime}\in\{\text{ Candidates }\}}\sum_{\text{ n-gram }^{\prime}\in C^{\prime}}\operatorname{Count}(\text{ n-gram }^{\prime})}.$ (20)
ROUGE-L score:
@@
$R_{\text{lcs}}=\frac{LCS(X,Y)}{m},$ (21)
$P_{\text{lcs}}=\frac{LCS(X,Y)}{n},$ (22)
$\mathbf{ROUGE-L}=\frac{(1+\beta^{2})R_{lcs}P_{lcs}}{R_{lcs}+\beta^{2}P_{lcs}}.$ (23)
METEOR score:
METEOR (Banerjee & Lavie, 2005) addresses certain inherent shortcomings of the BLEU score by taking into account both precision and recall evaluated over the entire corpus. The general formula for calculating METEOR score is as follows, where Penalty is the penalty of excessive word mismatches and $\alpha$ is a hyper-parameter:
$F=\frac{(\alpha^{2}+1)P}{R+\alpha P},$ (24)
$\mathbf{METEOR}=(1-\text{Penalty})\cdot F.$ (25)
BERT score:
BERT score (Zhang et al., 2019) is an automatic evaluation metric for text generation, which computes a similarity score for each token in the candidate sentence with each token in the reference sentence. The general formula for calculating BERT score is as follows, where tokens of reference sentence $x$ and candidate sentence $\hat{x}$ are represented by contextual embeddings:
$R_{\mathrm{BERT}}=\frac{1}{|x|}\sum_{x_{i}\in x}\max_{\hat{x}_{j}\in\hat{x}}\mathbf{x}_{i}^{\top}\mathbf{\hat{x}}_{j},$ (26)
$P_{\mathrm{BERT}}=\frac{1}{|\hat{x}|}\sum_{\hat{x}_{j}\in\hat{x}}\max_{x_{i}\in x}\mathbf{x}_{i}^{\top}\mathbf{\hat{x}}_{j},$ (27)
$F_{\mathrm{BERT}}=2\frac{P_{\mathrm{BERT}}\cdot R_{\mathrm{BERT}}}{P_{\mathrm{BERT}}+R_{\mathrm{BERT}}}.$ (28)
B.3 More performance comparison
Considering that there are additional related works capable of captioning input protein sequences but lacking instruction-following capabilities, we attempted to compare performance on overlapping properties and functions with these works. Table 12 shows a performance comparison with ProtT3 (Liu et al., 2024b), which is the most representative of these works. For open-ended generation and closed-set answer tasks, we selected Function, Similarity, Subcellular location, and Q&A related to structure and properties for comparison. The results demonstrate that our model exhibits significantly better performance.
Table 12: Performance comparisons on overlapping properties and functions.
B.4 More ablation studies
For the data ablation experiments, we also tested different forms of protein inputs, with the results shown in Table 13. For both forms of protein inputs, the addition of extra low-quality data does not result in performance improvement but instead led to performance degradation.
Table 13: Data ablation study on SEPIT’s pre-train data.

### v3

@@ Appendix B Construction of Protein Instruction Dataset
Currently, the widely recognized protein-text paired databases in the protein domain mainly include Swiss-Prot, TrEMBL (Bairoch and Apweiler, 1997), and RCSB PDB (Berman et al., 2000), with their specific contents detailed in Table 8. Since most of the text content in TrEMBL comes from automatic annotation methods, to eliminate the impact of its unreliability on the main experiments, unless otherwise specified, this paper only uses protein-text information from Swiss-Prot and RCSB PDB databases. The corresponding structures are from AlphafoldDB (Varadi et al., 2022) and experimentally determined structures in RCSB PDB. The statistical information about the complete protein instruction dataset is shown in Table 9 and Figure 4.
Table 8. Protein-text paired database.
@@
Table 9. Statistical information about the protein instruction dataset.
@@
Open-ended Generation Swiss-Prot Specific property or function 2,487,543 11,911
RCSB PDB Protein caption 125,703 2,500
Closed-set Answer RCSB PDB Specific property or function 1,131,327 22,500
GO-BP Biological process 601,381 24,162
GO-MF Molecular function 391,859 5,891
GO-CC Cellular component 327,780 6,735
EC Enzymatic catalytic activity 165,695 2,278
# All Instructions 5,231,288 75,977
@@
Figure 4. Distribution of protein sequence length and conversation length in the protein instruction dataset (training set).
@@
    • What is the primary function of ¡protein¿?
    • What is the main function of ¡protein¿?
    • What is the function of ¡protein¿?
    • Explain the function of ¡protein¿.
    • What is the characteristic function associated with the protein ¡protein¿?
    • Can you define the function profile of the ¡protein¿?
    • Give me the function caption of ¡protein¿.
      Example Response:
@@
    • Which protein family does ¡protein¿ belong to?
    • What is the protein family of ¡protein¿?
    • What is the closest related protein family for ¡protein¿?
    • Can you identify the family or group that ¡protein¿ belongs to?
    • To which protein family ¡protein¿ is classified?
    • Which protein class does ¡protein¿ fall into?
      Example Response:
@@
    • Where is ¡protein¿ located in the cell?
    • Can you specify the subcellular location of ¡protein¿?
    • What is the subcellular location of ¡protein¿?
    • Could you describe the subcellular location of ¡protein¿?
    • What are the primary subcellular regions where ¡protein¿ is detected?
      Example Response:
@@
    • Description the effects of environmental factors of ¡protein¿’s expression.
    • What are the environmental factors that induce the expression of ¡protein¿?
    • What environmental factors causes the upregulation of ¡protein¿?
    • What are the environmental factors that lead to the upregulation of ¡protein¿?
      Example Response:
@@
    • Which GO molecular function terms have ¡protein¿ been assigned to?
    • What molecular function is associated with ¡protein¿?
    • Which GO terms outline the functional capabilities of ¡protein¿?
    • What are the molecular functions of ¡protein¿?
      Example Response:
@@
    • Which GO biological process terms have ¡protein¿ been assigned to?
    • What biological process is associated with ¡protein¿?
    • Which GO terms outline the biological processes of ¡protein¿?
    • What biological processes is ¡protein¿ involved in, based on gene ontology annotations?
    • What are the biological processes of ¡protein¿?
      Example Response:
@@
    • Which GO cellular component terms have ¡protein¿ been assigned to?
    • What cellular component is associated with ¡protein¿?
    • Which GO terms outline the cellular components of ¡protein¿?
    • What cellular components is ¡protein¿ involved in, based on gene ontology annotations?
    • What are the cellular components of ¡protein¿?
      Example Response:
@@
    • At which specific developmental stages is ¡protein¿ expressed?
    • What are the developmental stages where ¡protein¿ is expressed?
    • What are the developmental stages where ¡protein¿ is detected?
    • What are the developmental stages where ¡protein¿ is found?
      Example Response:
@@
    • Can you identify and list all the motifs that are predicted to be present in ¡protein¿?
    • What are the short sequence motifs that are predicted to be present in ¡protein¿?
    • What are the short sequence motifs that are present in ¡protein¿?
    • What are the short sequence motifs that are found in ¡protein¿?
      Example Response:
@@
    • In which tissues is the expression of ¡protein¿ absent?
    • Describe the tissue-specific expression pattern of ¡protein¿?
    • What is the tissue-specific expression pattern of ¡protein¿?
    • What are the tissues where ¡protein¿ is expressed?
      Example Response:
@@
    • Describe the activity regulatory mechanism of ¡protein¿ associated enzymes, transporters, microbial transcription factors.
    • What is the activity regulatory mechanism of ¡protein¿?
    • Tell me about the activity regulatory mechanism of ¡protein¿.
      Example Response:
@@
    • What is the role of ¡protein¿ in the metabolic pathway?
    • Which metabolic pathway does ¡protein¿ associate with?
    • What is the metabolic pathway that ¡protein¿ is involved in?
      Example Response:
@@
    • Tell me about this protein ¡protein¿.
    • Give me some information about ¡protein¿.
    • Give me the abstract of ¡protein¿.
    • Give me a comprehensive description of ¡protein¿.
    • Tell me about ¡protein¿.
      Example Response:
@@
    • Does this protein contain non-polymer entities, ¡protein¿?
    • Does this protein contain polymer entities, ¡protein¿?
    • Does this protein contain DNA polymer entities, ¡protein¿?
    • Does this protein contain RNA polymer entities, ¡protein¿?
    • Does this protein contain solvent entities, ¡protein¿?
    • Does this protein contain branched entities, ¡protein¿?
    • Does this protein have unmodeled polymer monomers, ¡protein¿?
    • Does this protein have hybrid nucleic acid polymer entities, ¡protein¿?
    • Does this protein have cis-peptide linkages, ¡protein¿?
      Example Response:
@@
    • Does ¡protein¿ associate with enzyme classification ”{EC}”?
    • Does EC term ”{EC}” outline the enzyme classifications of ¡protein¿?
    • Is ¡protein¿ involved in enzyme classification ”{EC}”?
      Example Response:
@@
    • Does ¡protein¿ associate with biological process ”{GO}”?
    • Does GO term ”{GO}” outline the biological processes of ¡protein¿?
    • Is ¡protein¿ involved in biological process ”GO”?
      Example Response:
@@
    • Does ¡protein¿ associate with cellular component ”{GO}”?
    • Does GO term ”{GO}” outline the cellular components of ¡protein¿?
    • Is ¡protein¿ involved in cellular component ”{GO}”?
      Example Response:
@@
    • Does ¡protein¿ associate with molecular function ”{GO}”?
    • Does GO term ”{GO}” outline the functional capabilities of ¡protein¿?
    • Does ¡pro tein¿ have molecular function ”{GO}”?
      Example Response:
@@
Appendix C Supplement to the Experiments
C.1. Implementation Details
As depicted in Figure 2, we extensively utilize group learning rates throughout the entire training pipeline of SEPIT. We tend to assign higher learning rates to randomly initialized parameters, while opting for lower learning rates for pre-trained parameters in order to mitigate forgetting, setting the ratio between lower and higher learning rates at 0.1. In Stage 0, we actually employ ESM2-650M (Lin et al., 2023) as the pLM and PubMedBert (Gu et al., 2021) as the text encoder (Bio-BERT) to better encode biomedical text. We set the number of Gaussian Basis Kernels to 128. In Stage 1, we choose the representation from the penultimate layer of the protein encoder as input to LLMs to minimize the discrepancy between pre-training tasks and the current task. For the LLMs, we opt for TinyLlama-1.1B (Zhang et al., 2024). In Stage 2, we continue most of the settings from Stage 1, while setting the number of experts to 4, with Top-1 expert being activated at a time. At this stage, our protein encoder is frozen. Regarding the hyper-parameter settings for SEPIT-TinyLlama-MoEs, we set higher learning rate to $5e^{-5}$ and trained for 5 epochs in Stage 0. In Stage 1, we set higher learning rate to $2e^{-5}$ and trained for 1 epoch. In Stage 2, we set higher learning rate to $5e^{-5}$ and trained for 1 epoch. For all stages, we trained on 32 Tesla V100 GPUs, with a batch size per GPU set to 4, employing a warm-up and linear decay learning rate scheduler, and set the warm-up ratio to 0.06. In all experiments, we employ AMP, Zero Optimizer (Rasley et al., 2020) based on AdamW (Loshchilov and Hutter, 2019) and gradient checkpointing. For the hyper-parameter settings of other models, see Table 11. For all MoE models, we implement them based on DeepSpeed-MoE, employing expert parallelism during the training process and setting the expert parallel size to 4.
@@ C.1. Implementation Details
Table 10. Tokens consumed by API-based models.
@@
GPT-4o-mini $\sim$19M
GPT-4o $\sim$0.95M *
DeepSeek-V3 $\sim$19M
Table 11. Hyper-parameters of all models.
@@
Table 12. Training cost of all models.
@@
Table 13. Inference cost of main models.
@@
C.2. Metric Explanation
BLEU Score
The Bilingual Evaluation Understudy Score (Sutskever et al., 2014) (BLEU score) is a metric used to evaluate the quality of machine-translated text against human-translated reference texts, which is calculated using n-gram precision. The general formula for calculating BLEU score is as follows, where $\mathbf{BP}$ penalize overly short translations:
(16) $p_{n}=\frac{\sum_{C\in\{\text{ Candidates }\}}\sum_{\text{n-gram }\in C}\text{Count }_{\text{clip }}(\text{ n-gram })}{\sum_{C^{\prime}\in\{\text{ Candidates }\}}\sum_{\text{ n-gram }^{\prime}\in C^{\prime}}\operatorname{Count}\left(\text{ n-gram }^{\prime}\right)},$
(17) $\mathbf{BP}=\begin{cases}1&\text{if }c>r\\ e^{(1-r/c)}&\text{if }c\leq r\end{cases},$
(18) $\mathbf{BLEU}=\mathbf{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),$
(19) $\log\mathbf{BLEU}=\min(1-\frac{r}{c},0)+\sum_{n=1}^{N}w_{n}\log p_{n}.$
ROUGE-N Score
@@ C.2. Metric Explanation
(20) $\mathbf{ROUGE-N}=\frac{\sum_{C\in\{\text{ Candidates }\}}\sum_{\text{ n-gram }\in C}\operatorname{Count}_{\text{match}}(\text{ n-gram })}{\sum_{C^{\prime}\in\{\text{ Candidates }\}}\sum_{\text{ n-gram }^{\prime}\in C^{\prime}}\operatorname{Count}(\text{ n-gram }^{\prime})}.$
ROUGE-L Score
@@
(21) $R_{\text{lcs}}=\frac{LCS(X,Y)}{m},$
(22) $P_{\text{lcs}}=\frac{LCS(X,Y)}{n},$
(23) $\mathbf{ROUGE-L}=\frac{(1+\beta^{2})R_{lcs}P_{lcs}}{R_{lcs}+\beta^{2}P_{lcs}}.$
METEOR Score
METEOR (Banerjee and Lavie, 2005) addresses certain inherent shortcomings of the BLEU score by taking into account both precision and recall evaluated over the entire corpus. The general formula for calculating METEOR score is as follows, where Penalty is the penalty of excessive word mismatches and $\alpha$ is a hyper-parameter:
(24) $F=\frac{(\alpha^{2}+1)P}{R+\alpha P},$
(25) $\mathbf{METEOR}=(1-\text{Penalty})\cdot F.$
BERT Score
BERT score (Zhang* et al., 2020) is an automatic evaluation metric for text generation, which computes a similarity score for each token in the candidate sentence with each token in the reference sentence. The general formula for calculating BERT score is as follows, where tokens of reference sentence $x$ and candidate sentence $\hat{x}$ are represented by contextual embeddings:
(26) $R_{\mathrm{BERT}}=\frac{1}{|x|}\sum_{x_{i}\in x}\max_{\hat{x}_{j}\in\hat{x}}\mathbf{x}_{i}^{\top}\mathbf{\hat{x}}_{j},$
(27) $P_{\mathrm{BERT}}=\frac{1}{|\hat{x}|}\sum_{\hat{x}_{j}\in\hat{x}}\max_{x_{i}\in x}\mathbf{x}_{i}^{\top}\mathbf{\hat{x}}_{j},$
(28) $F_{\mathrm{BERT}}=2\frac{P_{\mathrm{BERT}}\cdot R_{\mathrm{BERT}}}{P_{\mathrm{BERT}}+R_{\mathrm{BERT}}}.$
C.3. More Ablation Studies
For the data ablation experiments, we also tested different forms of protein inputs, with the results shown in Table 14. For both forms of protein inputs, the addition of extra low-quality data does not result in performance improvement but leads to performance degradation.
Table 14. Data ablation study on SEPIT’s pre-train data.